OpenStudy (anonymous):

find dy/dx at (4,pi/2) if r=2+3sin theta

OpenStudy (anonymous):

\[\frac{ dy }{ dx }at (4,\frac{ \pi }{2 }),, and,, r=2+3\sin \theta \]

OpenStudy (alekos):

are you looking for dy/dx or dr/d(theta) ?

OpenStudy (anonymous):

dy/dx

OpenStudy (alekos):

then this is a tricky problem, have you tried it?

OpenStudy (alekos):

OK. You need to use the following.... For r = f(θ) and x=rsinθ y=rcosθ (polar co-ordinates) dy/dx = [(dr/dθ)sinθ +rcosθ]/[(dr/dθ)cosθ - rsinθ)] You just need to find the derivative and plug in the co-ordinate point

OpenStudy (alekos):

If you're wondering how i got the fancy dy/dx formula i can send you the derivation

OpenStudy (alekos):

By the way the graph doesnt go through (4, pi/2), is this point correct?